Tom has 117 quarters and dimes worth a total of 15.75. How much of each does he have?? Thanks!!
Answer by checkley77(12844)

Q+D=117 or Q=117-D
.2RQ+.10D=15.75
.25(117-D)+.10D=15.75
29.25-.25D+.10D=15.75
-15D=15.75-29.25
-.15D=-13.50
D=-13.50/-.15
D=90 DIMES.
Q=117-90
Q=27 QUARTERS.
PROOF:
.25*27+90*.10=15.75
6.75+9.00=15.75
15.75=15.75